Queens Award for Voluntary Service

FFOG Minutes: 14 March 2024

  • Location: Friends Meeting House Rm 4 (On-site meeting relocated to FMH room 4 due to bad weather.)
  • Date: 14.03.24
  • Time: 2.00pm to 4.00pm


Sue Nieduszynska (chair), Graham Brandwood Steve Bullen, David Elliot, Clio Graham, Anne Greenwood, Rebekah Mulligan, Ian Procter, Hilary Short, Mick Short, Hilary Smith, Chris Workman.


David Brookes.

Previous Minutes


Matters Arising

Ear tags

In Iiem 7.5 in the previous minutes it was documented that Robin had stated that ear tags are only a legal requirement when cattle are being moved from one establishment to another.GB said that this was incorrect and requested that this should be recorded. SN said that ear tags were the responsibility of the graziers and that any further discussion on the topic should take place outside of the FFOG meeting..

Agenda Items

2.1.2. Feedback from the FA. 

Legacy. HS stated that a large sum of money has been left to the FA & there will be a meeting next Tuesday to discuss how it would be appropriate to use this legacy. Currently, the preferred option is to purchase the field of approx. 2 acres adjacent to Pony Wood just before the Canal. Two unsuccessful approaches have been made via the Land Agent.  SN has managed to track down a friend of the people on Guernsey and it is hoped that a personal touch might prove more successful.

Other possible uses could be education:

  • Tony has proposed a dipping pond. Suggestions: in the wet area close to the Orchard Footpath, or the Paddock or building a pier near School Pond.
  • Improvements to the stump circle area.
  • Outreach education.
  • Improvements to and ponds near Lucy Brook.

New Website. SN reported that Mark Tanner is continuing to work on this. When there is new information or there are new additions to the blog, this will be displayed in a window on the home page.SN wishes to encourage those who spend a lot of time on the Reserve to contribute to the blog. Currently, material for the blog is emailed to Mark (e.g. Via “Report a Sighting” on the homepage. Another feature will be a link to species identification websites. SN enquired if there was any interest in involvement in species identification (other than birds).                 

2.1.3. New Saplings for 2025

RM said a group had identified 90 places for sapling planting in Pony Wood. In addition, IP said he had 3 other sites in mind. A group had identified 90 places for sapling planting in Pony Wood. In addition, IP said he had 3 other sites in mind. RM to convene a subgroup to look at potential sites for sapling planting.

2.2. Orchard Willow Domes

The Orchard Willow domes at the eastern end have not taken because of shade from the Ash tree. GB proposed a rebuild using a mixture of Hazel & Willow. SN proposed “stepping stumps/logs”. It was agreed that there is space for both proposals.

2.3. Snipe Flush

See report from CG. Numbers were a little bit down this year. However, the numbers are variable dependent on local weather conditions. Dan Haywood has indicated that numbers are generally declining in the area but that there is no hard data.SN suggested that our snipe data, currently stored on drop box should go on the website along with other surveys & that we should also include the rainfall figures for Lancaster over years in the flush count report. (information obtained from Lancaster University) It was also discussed that the weather conditions in Northern & Eastern Europe could influence migration & the snipe counts. Local weather conditions may also influence the relative numbers of snipe in the different areas of the Reserve. Dropbox will be discontinued and replaced by Google Workspace.SN to inform Glenys about changes to drop box.

2.4. New Hedge in West Field

Saplings were planted and an electric fence erected. Unfortunately, the cattle walked straight through the fence. Either it was faulty, or the older cattle have learned that it is possible to walk through it as they have had experience of fences that were not switched on. They were probably heading for the shelter of the hedge. The fence has been rebuilt and it has been tested. IP suggested a wire fence should be put in place at a cost of about £300. It was agreed that a protective wire fence would be erected in West Field. RM will order posts & organise a volunteer task force to install the fence. HS will discuss the fence with Fraser and request he defer putting the cattle in until erected.

2.5. Long Pads. Gaps in the hedge allowing dogs to access Flora Field.

RM reported that approx. 10m stretch of hedge has been laid where the gaps were and brash used to fill other gaps.

2.6 Pony Wood and Cromwell Road gates.

See report RM. A member of the public was concerned about dogs hurting themselves pushing up against the wire mesh on the lower part of the Pony Wood fence. To rectify the issue with wire mesh under Pony Wood gate. RM to add to volunteer tasks. There is also a problem with Cromwell Rd gate. The current rope fastening, which although quiet, is difficult for blind/partially sighted people to use. Fraser had commented that the gates at the abattoir have fastenings that are metal encased in plastic.SN to investigate the gate fastening at the abattoir. CW to send details of an online agricultural supplier of gate materials.

2.7. Upper Sowerholme-Ensuring the boundary is secure for cattle.

SN reported that the cattle who had got into US had accessed it via the pedestrian entrance.GB has fixed the breach.

2.8. Fauna Path.

IP discussed trimming the hedge between the path & School Field.This refers to the hedge from school pond heading in the direction of Cromwell Rd. Robin looks after the hedge on his side and the top. It is growing closer to the path on our side. IP proposed that we cut it back by about 2 feet next winter. To inspect this hedge when the subgroup meets to discuss the siting of new saplings. RM to add this to the agenda for the subgroup meeting.

2.9. Hay & Carr House Meadows- Poaching & related issues.

At the outset SN said that there was a consensus from the trustees that existing agreements must be honoured. GB referred to the previous minutes (item 6.1.6) where a difference of opinion was recorded regarding the severity of poaching in Carr House Meadow. GB said that if he flags up an issue of concern in his Monitoring inspection, then it should be given due credence and not be dismissed as alarmist or not an issue.

SN said the monitoring inspections are reports for FFOG and that members of the group are at liberty to comment on the content. We also have guidance from NE and HS added that in the past we have submitted photographic evidence of “poached” areas to NE, on which we base our assessment. HS had subsequently spoken to Fraser who removed the cattle from the field although they have subsequently been in Hay Meadow.

IP said there was not an issue if the cattle were put into Hay Meadow for 5 to 10 days in September to graze off and he thought this had previously been agreed.  Hils was concerned that the current grazing pattern would bring unwanted weeds & nutrition into the Hay Meadow. She said the cattle should be put in to graze until the vegetation is 2cm to 10 cm and then not go back in for the rest of the year. Hils said that unless this was the case, there had to be acceptance that Hay meadow would never become a wildflower meadow and there was no point paying for surveys to be done.

MS stressed the importance of external monitoring. SN thought this was a separate issue and pointed out that we have to work within the current agreements. We do not control the graziers, but we negotiate with them. SN pointed out that the paper submitted by GB made no reference to soil pH which may be a significant factor. HS indicated that it might be possible in a few years’ time to have one single grazing episode per annum but could not commit until we enter a new agreement. HS & MS said we mustn’t lose sight of the improvement in the Meadow to date, although accepted that it had now levelled off.

GB proposed an experiment to install an electric fence in Hay Meadow 10m in and running parallel to fauna path to see if there is a difference and an improvement in wildflowers. IP said it would take several years to see a change. IP added that there was an existing box for the electric fence and that we could use to restrict cattle access to a section near the double gates. HS suggested a strip that ran at 90 degrees to the fauna path. There were differing views on the management of the Hay Meadow and the level of cooperation/negotiation with the graziers. Trustee agreement to continue with independent surveys. This gives long term consistency over the duration the NE agreement. To consider GB’s experiment in Hay Meadow at the next site meeting May 2nd.

2.10 Willow Screen

SN queried the willow management plan. The minutes of July ’22 FFOG were different from those of the last FFOG meeting. GB said that the willow screen was an important habitat in its own right although its purpose initially was to protect birds on school pond. The channel between the screen and the brash is a bird habitat, there is extensive growth of moss and lichens on the willow. The willow is also farmed. SN had understood that the willow at the northern end would be allowed to mature and produce catkins & foraging for insects in the spring. GB said the willow at the northern end was allowed to produce catkins last year by delaying cutting until summer. IP said approximately half is cut each year and this allows both catkin development and willow to grow to a size that can be farmed. SN queried whether we should allow some develop. SN suggested to inspect the area at the May site visit.

3.1. Organised Events

  • Monday 1 April -Easter Egg Hunt
  • Sunday 16 June- NLWG Wildlife Fair

3.2. Financial Control.

Nothing to report.


Arrangements & person responsible.AG suggested that for each arrangement there should be and identified person responsible. AG to suggest the responsible person when the arrangements are next sent out. FFOG members to feedback if this needs to be altered. Arrangements for April. No 5. IP suggested we remove ‘annual review of herbicide spraying COSHH & site assessment forms due”.  Herbicide has not been used for about 8 years. This was agreed.                                                                                                                                            

GB reported that the bull is now back with the herd.

Next meeting

Onsite 2 May 2024 5.00pm to 7.00pm

Download PDF

PDF of Decisions and Actions 14 March 2024